OpenStudy (anonymous):

What would y be when x=-0.8

OpenStudy (lncognlto):

What is the equation?

OpenStudy (anonymous):

yeah, you dont have an equation :/

OpenStudy (anonymous):

y=3x+1

OpenStudy (lncognlto):

Just substitute the value of x into the equation y = 3(-0.8) + 1 =?

OpenStudy (anonymous):

substitute your x with its value , then solve

OpenStudy (anonymous):

I dont get it?

OpenStudy (lncognlto):

Replace x in the equation with the x = value in the question.

OpenStudy (lncognlto):

y = 3x + 1 and x = -0.8 so y = 3(-0.8) + 1

OpenStudy (anonymous):

haha ow come on, Incognlto already did it for you! just need to solve it :) coz we're teaching you how to fish, & not giving the fish to you directly :)

OpenStudy (anonymous):

I get it now thanks

OpenStudy (lncognlto):

You're welcome :)

OpenStudy (anonymous):

the value of x when y = 8.2?

OpenStudy (anonymous):

Solve the inequality 4x-3<7?

OpenStudy (lncognlto):

Now you need to solve for x, so put x on one side and all the other values on the other. What do you get?

OpenStudy (lncognlto):

That works for both, but lets just do the first one now.

OpenStudy (lncognlto):

Come on, 8.2 = 3(x) + 1... What next?